# Geometric QCD II: The Confining Twistor String and Meson Spectrum

### 12 Mar 2026 v3

@@ Appendix E The Pinch Singularity and the Topological Barrier
$I(z_{0})=\frac{-\mathrm{i}}{2}\int_{z\bar{z}<1}dzd\bar{z}\,\partial_{\bar{z}}\left(\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}\right)$ (E.2)
@@
By applying Stokes’ theorem, the two-dimensional bulk integral over the disk $D$ reduces exactly to a one-dimensional contour integral over the boundary unit circle $z\bar{z}=1$:
$I(z_{0})=\frac{-\mathrm{i}}{2}\int_{z\bar{z}=1}dz\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}$ (E.3)
@@
Crucially, on the unit circle, the complex conjugate is strictly given by the geometric inversion $\bar{z}=1/z$. Substituting this into the logarithm yields the exact boundary representation of the action:
$I(z_{0})=\frac{-\mathrm{i}}{2}\oint_{|z|=1}dz\frac{\log(1/z-\bar{z}_{0})}{z-z_{0}}$ (E.4)
This exact formula explicitly reveals the analytic structure governing the boundary dynamics. The integrand possesses a simple pole at $z=z_{0}$ (originating from the twistor itself) and a logarithmic branch point at $z=1/\bar{z}_{0}$ (originating from the boundary inversion of the conjugate twistor).
At $z_{0}=0$ this integral reduces to
$I(0)=\frac{-\mathrm{i}}{2}\int_{0}^{2\pi}\mathrm{i}\,d\theta\left(-\mathrm{i}\theta\right)=\frac{-\mathrm{i}}{2}\frac{(2\pi)^{2}}{2}=-\mathrm{i}\pi^{2}$ (E.5)
which matches with the computation of the Minkowski-continued area in the previous section.
@@
Because the physical domain restricts the pole to the interior of the disk ($|z_{0}|<1$), its conjugate mirror strictly resides outside the disk ($|1/\bar{z}_{0}|>1$). However, if any continuous Langevin fluctuation attempts to push the twistor pole across the boundary ($|z_{0}|\to 1$), its conjugate mirror must simultaneously approach the boundary from the outside.
@@
At the exact moment of crossing, these two singularities violently collide directly on the physical integration contour $|z|=1$, trapping the integration path between them. This creates a severe, non-integrable pinch singularity, causing the action integral to macroscopically diverge. This geometric divergence acts as an infinite energetic barrier, rigorously protecting the twistor pole from escaping and locking the string into its quantized topological sector.
Appendix F Algebraic Derivation of Spin Holonomy and Trace T
To rigorously justify the topological selection rules governing the meson spectrum, we compute the path-ordered Dirac trace $T$ for the one-pole sector ($z_{0}=0$). The fundamental link matrix $\mathcal{M}(\theta)$ at the boundary $|z|=1$ is defined as:
$\mathcal{M}(\theta)=i\mathcal{P}(\theta)=-2\left(e^{i\theta}\eta\xi^{\dagger}-e^{-i\theta}\xi\eta^{\dagger}\right)$ (F.1)
where $\eta$ is the basis spinor of the frozen $SU(2)_{R}$ sector and $\xi(\theta)$ is the rotating spinor of the active $SU(2)_{L}$ sector. In the basis where $\eta=\begin{pmatrix}1\\ 0\end{pmatrix}$ and the twistor $\lambda(z)\propto 1/z$ implies $\xi(\theta)=\begin{pmatrix}0\\ e^{-i\theta}\end{pmatrix}$ at the boundary, the local matrix evaluates to:
$\mathcal{M}(\theta)=\begin{pmatrix}0&-2e^{2i\theta}\\ 2e^{-2i\theta}&0\end{pmatrix}$ (F.2)
This matrix represents a local $SU(2)$ rotation that flips the chiral state. The propagator $\mathcal{U}$ for one physical cycle is the path-ordered product of these matrices. For two adjacent steps $\theta_{1}$ and $\theta_{2}=\theta_{1}+\Delta\theta$, we find:
$\mathcal{M}(\theta_{2})\mathcal{M}(\theta_{1})=\begin{pmatrix}-4e^{2i\Delta\theta}&0\\ 0&-4e^{-2i\Delta\theta}\end{pmatrix}$ (F.3)
Assuming even number $2L$ of points on a unit circle in discretized loop, and multiplying these matrices around the entire loop $\oint d\theta=2\pi$ reduces the propagator for one cycle to:
$\mathcal{U}=(-4)^{L}\begin{pmatrix}e^{2\pi i}&0\\ 0&e^{-2\pi i}\end{pmatrix}$ (F.4)
The physical trace $T$
[…]
The real factor renormalizes the bare mass of quark, so we drop it, but the phase is interesting, as ir accumulates in the windings and leads to the negative sign in front of $\exp\left(\Delta S\right)$, which changes the quantization condition by selecting odd or even n in B-S quantization $\Delta S_{eff}=\pi n$ depending on the parity of $L$.
Appendix G Geometric Potentials and Universal Trajectories
In this appendix, we decouple the unknown dimensionful string tension $\sigma$ from the geometric quantization conditions to derive the parametric functions $h(a)$ and $g(a)$, their singularities, and their roots.
We formulate the system directly in terms of the fractional winding parameter $a$. As derived in Section 15.5, the macroscopic spatial rotation angle of the string is exactly $\phi=4\pi a$. We isolate the string tension from the boundary potential $F(a)$ by defining the dimensionless geometric function:
$\tilde{F}(a)=\frac{2\sin^{4}(\pi a)}{1-a}$ (G.1)
The boundary potential is $F(a)=\frac{1}{\pi\sigma}\tilde{F}(a)$. Defining the physical mass-squared as $m^{2}=E^{2}=\sigma h(a)$, the unknown tension $\sigma$ factors to the numerator of the mass equation and cancels out of the angular momentum equation entirely.
The total geometric phase for the $q$-shifted trace sector is:
$\Phi_{q}(a)=-E^{2}F(a)+\frac{\pi a^{2}}{12}+4\pi aJ+2\pi qa$ (G.2)
where $q=0$ for the unshifted $\pi/K$ integral, and q=1q=-1 for the shifted $\rho$ and Scalar integrals.
We impose the saddle-point condition $\Phi_{q}^{\prime}(a)=0$ to isolate the angular momentum terms:
$4\pi aJ+2\pi qa=-\frac{\pi a^{2}}{6}+aE^{2}F^{\prime}(a)$ (G.3)
Substituting this into the Bohr-Sommerfeld quantization condition $\Phi_{q}(a)=2\pi n_{w}$ (where $n_{w}=0$ for the principal string sector), we obtain the algebraic relation:
$-\frac{\pi a^{2}}{12}-E^{2}F(a)+aE^{2}F^{\prime}(a)=0$ (G.4)
Using the differential identity $aF^{\prime}(a)-F(a)=\frac{1}{\pi\sigma}\left(a\tilde{F}^{\prime}(a)-\tilde{F}(a)\right)$, we isolate $m^{2}=\sigma h(a)$ to obtain the mass function:
$h(a)=\frac{a^{2}\pi^{2}}{12\left(a\tilde{F}^{\prime}(a)-\tilde{F}(a)\right)}$ (G.5)
Simplifying that expression, we get an explicit trigonometric answer:
$h(a)=-\frac{\pi^{2}(a-1)^{2}a^{2}\csc^{4}(\pi a)}{24(-2a+4\pi(a-1)a\cot(\pi a)+1)}$ (G.6)
To obtain the angular momentum function, we substitute $E^{2}=\sigma h(a)$ back into the saddle-point equation. Defining the target angular momentum as $u=J+q/2$, we obtain:
$u=\frac{a\left(a\tilde{F}^{\prime}(a)-2\tilde{F}(a)\right)}{48\left(\tilde{F}(a)-a\tilde{F}^{\prime}(a)\right)}$ (G.7)
This expression can be dramatically simplified. Dividing the numerator and denominator by $\tilde{F}(a)$ to isolate the logarithmic derivative $\tilde{F}^{\prime}(a)/\tilde{F}(a)$, we obtain a partial fraction decomposition:
$\frac{\tilde{F}^{\prime}(a)}{\tilde{F}(a)}=\frac{2a+48u}{a(a+48u)}=\frac{1}{a}+\frac{1}{a+48u}$ (G.8)
Substituting the explicit logarithmic derivative $\tilde{F}^{\prime}(a)/\tilde{F}(a)=4\pi\cot(\pi a)+\frac{1}{1-a}$ and combining the kinematic poles yields the exact trigonometric spectral equation:
$4\pi\cot(\pi a)=\frac{1-2a}{a(1-a)}+\frac{1}{a+24(2J+q)}$ (G.9)
The Regge trajectory can also be simplified to the following
$J+\frac{q}{2}=\frac{\sin^{4}(\pi a)}{2\pi^{2}a(1-a)}\frac{m^{2}}{\sigma}-\frac{a}{48}$ (G.10)
G.1 Singularities and the Ground States
The geometry of the functions is dominated by a kinematic pole at the speed-of-light boundary, $a\rightarrow 1/2$. At this limit, the left-hand side of (G.9) strictly vanishes ($\cot(\pi/2)=0$). For the equation to hold, the right-hand side must also vanish, which dynamically forces the angular momentum to diverge ($u\rightarrow\infty$) and forms the linear macroscopic Regge asymptote. Evaluating the derivative $\tilde{F}^{\prime}(1/2)=8$ in the mass formula (G.5) restores the macroscopic Regge slope $J\approx 2m^{2}/(\pi^{2}\sigma)$. The asymptotic expansion at larger $s=m^{2}/\sigma$ corresponds to an inverted expansion around $a\to 1/2-\epsilon$. The expansion of trajectory goes as follows:
$J(s)$ $=\frac{2s}{\pi^{2}}+B_{0}+\frac{B_{1}}{s}+\frac{B_{3}}{s^{3}}+O(s^{-5});$ (G.11)
$B_{0}$ $=-\left(\frac{q}{2}+\frac{1}{96}\right);$ (G.12)
$B_{1}$ $=\frac{\pi^{2}}{36864\left(\pi^{2}-2\right)}$ (G.13)
$B_{3}$ $=\frac{\pi^{6}\left(48-24\pi^{2}+5\pi^{4}\right)}{32614907904\left(\pi^{2}-2\right)^{4}}$ (G.14)
The coefficients $B_{1},B_{3}$ are numerically so small, that the trajectory is indistinguishable from linear for $J>1$.
When we compute the roots for the ground states, the explicit spectral equation (G.9) yields the mass eigenvalues:
  • 1.
    The Light Pseudoscalar ($q=0,u=0$): Solving (G.9) yields $a_{\pi}\approx 0.4268$. Substituting into $h(a)$ gives the mass ratio $m_{\pi}^{2}/\sigma\approx 0.0375$. The geometry produces a light ground state without symmetry breaking parameters.
  • 2.
    The Vector Ground State (q=1,u=1/2q=-1,u=1/2): The topological shift forces the string to spin faster. Solving (G.9) occurs at $a_{s}\approx 0.4987$, yielding $m_{s}^{2}/\sigma\approx 1.9782$. The geometry predicts the vector state to be massive.
The continuous generation of the parametric trajectories $u=J+q/2$, the numerical roots, and the data fits to $\sigma$ are performed using the companion Mathematica notebook [35].
G.2 The Endpoint Singularity and the Fold Catastrophe
While the large-$s$ asymptotic expansion (G.11) governs the strictly linear macroscopic Regge regime, the exact parametric geometry of the twistor string dictates a hard geometric boundary at small $s$. This boundary manifests not as a smooth continuation toward $s\to$-$\infty$, but as a fold catastrophe in the mapping between the spectral parameter $a$ and the physical observables.
To expose this geometric termination, we rewrite the exact parametric equations for the mass ratio $s(a)=h(a)=m^{2}/\sigma$ and the target angular momentum $u(a)=J+q/2$. From (G.6) and (G.8), we define the auxiliary functions:
$\mathcal{P}(a)$ $=\frac{\pi^{2}a(1-a)}{24\sin^{4}(\pi a)},$ (G.15)
$\mathcal{R}(a)$ $=4\pi\cot(\pi a)-\frac{1-2a}{a(1-a)}.$ (G.16)
The spectral equations then take the remarkably compact algebraic form:
$s(a)$ $=\mathcal{P}(a)\mathcal{R}(a)^{-1},$ (G.17)
$48u(a)$ $=\mathcal{R}(a)^{-1}-a.$ (G.18)
A fundamental property of the Rigid Twistor String is the existence of a simultaneous mathematical endpoint $a_{c}$ where the first derivatives of both the mass-squared and the angular momentum vanish exactly. We can rigorously prove this simultaneous singularity by differentiating the target angular momentum:
$48u^{\prime}(a)$ $=-\mathcal{R}^{\prime}(a)\mathcal{R}(a)^{-2}-1=0$
$\implies-\mathcal{R}^{\prime}(a)=\mathcal{R}(a)^{2}.$ (G.19)
Next, we differentiate the mass function and multiply by $\mathcal{R}(a)^{2}$:
$s^{\prime}(a)\mathcal{R}(a)^{2}=\mathcal{R}(a)\left[\mathcal{P}^{\prime}(a)+\mathcal{P}(a)\mathcal{R}(a)\right].$ (G.20)
By explicitly taking the logarithmic derivative of $\mathcal{P}(a)$, we uncover an exact geometric identity:
$\frac{\mathcal{P}^{\prime}(a)}{\mathcal{P}(a)}=\frac{1-2a}{a(1-a)}-4\pi\cot(\pi a)\equiv-\mathcal{R}(a).$ (G.21)
Consequently, the term in the brackets vanishes identically. Thus, we arrive at the exact algebraic statement: $s^{\prime}(a_{c})=0$ precisely whenever $u^{\prime}(a_{c})=0$.
In the physical $(s,J)$ plane, this simultaneous extremum signifies a fold catastrophe. The parametric Regge trajectory reaches a hard geometric endpoint where the physical slope $dJ/ds=u^{\prime}(a)/s^{\prime}(a)$ resolves (via L’Hôpital’s rule) to a finite limit $S_{0}$, while the trajectory curvature diverges to infinity, $J^{\prime\prime}(s)\to\infty$.
The critical parameter $a_{c}$ is determined by the exact transcendental equation $\mathcal{R}^{\prime}(a)=\mathcal{R}(a)^{2}$, which evaluates to:
$4\pi^{2}\csc^{2}(\pi a)-\frac{1-2a+2a^{2}}{a^{2}(1-a)^{2}}$
$=\left(4\pi\cot(\pi a)-\frac{1-2a}{a(1-a)}\right)^{2}.$ (G.22)
Solving this within the fundamental domain $a\in(0,1/2)$ yields the universal critical parameters:
Crit. Param. ($a_{c}$): $\approx 0.311277$
Thres. Mass Rat. ($s_{c}=m_{c}^{2}/\sigma$): $\approx 0.027791$
Crit. $\pi$ Spin ($q=0$): $J_{c}\approx-0.003378$
Crit. $\rho$ Spin (q=1q=-1): $J_{c}\approx 0.496622$ (G.23)
Figure 16: The exact geometric Regge trajectories plotted down to the mathematical endpoint $(s_{c},J_{c})$. The trajectory departs from this boundary with a finite slope ($S_{0}\approx 0.1118$) that is smaller than the asymptotic slope at large $s$ ($\approx 2/\pi^{2}\approx 0.2026$). The infinite curvature $J^{\prime\prime}(s)\to\infty$ at the exact endpoint is visually imperceptible. To the eye, this looks like a perfectly linear trajectory ending abruptly at this point.
At this mathematical endpoint, the Regge trajectory departs with a finite slope $S_{0}\approx 0.1118$ and infinite curvature (see Fig. 16). While this fold catastrophe represents a severe singularity in the complexified string geometry, it is essentially invisible to the naked eye on the physical Chew-Frautschi plot (Figure 15). The physical branch of the spectrum strictly corresponds to $u\geq 0$. As established previously, the physical ground state terminates at $a_{min}$ where $u=0$. For the unshifted pion trajectory ($q=0$), this occurs at $a_{min}\approx 0.4268$.
Because the physical root $a_{min}$ sits comfortably above the critical parameter $a_{c}$ ($0.4268>0.3113$), the trajectory mathematically terminates well before the string can experience the fold catastrophe. This mass gap ($s(a_{min})>s_{c}>0$) acts as an impenetrable geometric shield. It guarantees that the Regge trajectory never crosses into the $s<0$ regime, rigorously protecting the Planar QCD spectrum from tachyonic instabilities while enforcing the exact reality of all physical eigenvalues. The phenomenological Regge “intercepts” $\alpha_{\pi,\rho}(0)$ are thus not true physical states, but simply linear extrapolations of the large-$s$ asymptotic expansion bridging over this hidden, non-linear geometric boundary.
While this fold catastrophe represents a severe singularity in the complexified string geometry, it is essentially invisible to the naked eye on the physical Chew-Frautschi plot (Figure 15). The physical branch of the spectrum strictly corresponds to $u\geq 0$. As established previously, the physical ground state terminates at $a_{min}$ where $u=0$. For the unshifted pion trajectory ($q=0$), this occurs at $a_{min}\approx 0.4268$.
Appendix H Irrationality of the Physical Branch Points
As discussed in Section 15.8, rational values of the branch point degree $a=m/n$ would cause the twistor trajectory to close after finitely many windings. In that case, all distinct states would be counted with just $2n$ windings, and our mechanism for the Exact WKB spectrum would disappear. In this appendix, we rigorously prove that the physical spectrum is mathematically protected from these rational degeneracies.
Theorem. For any rational target angular momentum $u=J+q/2\geq 0$, the exact spectral equation (G.9) possesses no rational solutions for $a\in(0,1/2)$.
Proof. The exact geometric quantization is governed by the spectral equation derived in Appendix G:
$4\pi\cot(\pi a)=\frac{1-2a}{a(1-a)}+\frac{1}{a+48u}$ (H.1)
Assume, for the sake of contradiction, that the physical branch point $a$ is a rational number $m/n\in(0,1/2)$. Since the target momentum $u$ is also rational by definition, the algebraic expression on the right-hand side evaluates to a purely rational number.
On the left-hand side, it is a foundational property of trigonometric functions that for any rational argument $a$, the value of $\cot(\pi a)$ is an algebraic number. Furthermore, for $a\in(0,1/2)$, the value $\cot(\pi a)$ is strictly positive and therefore non-zero. By the Lindemann–Weierstrass theorem, $\pi$ is a transcendental number. The product of a transcendental number and a non-zero algebraic number is strictly transcendental.
Consequently, the left-hand side is a transcendental number, while the right-hand side is a rational number. This is a strict mathematical contradiction. Therefore, the branch point parameter $a$ cannot be a rational number. $\blacksquare$

### 23 Mar 2026 v4

@@ Appendix D The Pinch Singularity and the Topological Barrier
$I(z_{0})=\frac{-\mathrm{i}}{2}\int_{z\bar{z}<1}dzd\bar{z}\,\partial_{\bar{z}}\left(\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}\right)$ (D.2)
@@
By applying Stokes’ theorem, the two-dimensional bulk integral over the disk $D$ reduces to a one-dimensional contour integral over the boundary unit circle $z\bar{z}=1$:
$I(z_{0})=\frac{-\mathrm{i}}{2}\int_{z\bar{z}=1}dz\frac{\log(\bar{z}-\bar{z}_{0})}{z-z_{0}}$ (D.3)
@@
Crucially, on the unit circle, the complex conjugate is given by the geometric inversion $\bar{z}=1/z$. Substituting this into the logarithm yields the boundary representation of the action:
$I(z_{0})=\frac{-\mathrm{i}}{2}\oint_{|z|=1}dz\frac{\log(1/z-\bar{z}_{0})}{z-z_{0}}$ (D.4)
@@
This formula explicitly reveals the analytic structure governing the boundary dynamics. The integrand possesses a simple pole at $z=z_{0}$ (originating from the twistor itself) and a logarithmic branch point at $z=1/\bar{z}_{0}$ (originating from the boundary inversion of the conjugate twistor). Because the physical domain restricts the pole to the interior of the disk ($|z_{0}|<1$), its conjugate mirror resides outside the disk ($|1/\bar{z}_{0}|>1$). However, if any continuous Langevin fluctuation attempts to push the twistor pole across the boundary ($|z_{0}|\to 1$), its conjugate mirror must simultaneously approach the boundary from the outside.
@@
At the moment of crossing, these two singularities collide directly on the physical integration contour $|z|=1$, trapping the integration path between them. This creates a severe, non-integrable pinch singularity, causing the action integral to macroscopically diverge. This geometric divergence acts as an infinite energetic barrier, protecting the twistor pole from escaping and locking the string into its quantized topological sector.
Appendix E Analytical and numerical study of the Twistor String equation
The physical states of the twistor string are governed by the total action monodromy $\Delta S$ accumulated over one period of rotation. The Bohr-Sommerfeld quantization condition requires $\Delta S=2\pi n$. As derived in Section 16, factoring out the overall twistor scale parameter $a$, we can write $\Delta S=a\Phi$, where the phase functional is
$\Phi(K,\beta)$ $=2\pi EK-4\pi\left(J+\frac{q}{2}\right)-\pi\sigma K^{2}\left(\beta+\frac{\sin 2\beta}{2}\right)$
$-4\pi mK\cos\beta-\frac{2}{3}(\tan\beta-\beta).$ (E.1)
Here, $K=R/a$ is the kinematic radius, and $q$ represents the topological vector index ($q=0$ for the pseudo-scalar families, and q=1q=-1 for the vector families).
E.1 Significant reduction of the spectrum: only the $n=0$ level is left
The classical string trajectory is determined by the saddle-point equations with respect to the dynamical variables $a,K$, and $\beta$. Setting $\partial_{a}\Delta S=0$, we get the first equation:
$\Phi(K,\beta)=0,$ (E.2)
which implies that there are no levels with $n\neq 0$ in the general Bohr-Sommerfeld quantization condition:
$\Delta S=2\pi n,\qquad n=0,\pm 1,\pm 2,\dots$ (E.3)
This important conclusion follows from the fact that our $\Delta S$ depends linearly on $a$ at fixed $\beta$ and $K=R/a$. Therefore, by changing our variables to $a,K,\beta$, we completely eliminate $a$ from the saddle-point equations and also obtain the extra equation (E.2) for $\beta$ and $K$.
Next, setting $\partial_{K}\Phi=0$ yields the energy equation:
$E=\sigma K\left(\beta+\frac{\sin 2\beta}{2}\right)+2m\cos\beta.$ (E.4)
Setting the boundary variation $\partial_{\beta}\Phi=0$ balances the boundary tensions and yields a quadratic equation for $K$:
$\sigma\cos^{2}\beta\,K^{2}-2m\sin\beta\,K+\frac{1}{3\pi}\tan^{2}\beta=0.$ (E.5)
Note that we have obtained three equations for three unknowns $E,K,\beta$, with $a$ decoupled. Thus, we found the two equations for the two parameters of the saddle point and the third equation for the energy spectrum, just as intended, with the important restriction that all the extra Bohr-Sommerfeld levels $n\neq 0$ are absent.
E.2 Algebraic simplifications of the spectral equation
To systematically analyze and numerically fit these equations, it is advantageous to cast them into a dimensionless form. We introduce the dimensionless renormalized boundary mass parameters $x=m/\sqrt{\sigma}$ for the light quarks and $x^{\prime}=m^{\prime}/\sqrt{\sigma}$ for the strange trajectories, where $m^{\prime}=(m+m_{s})/2$. Solving the quadratic equation for the radius, we find the analytical root:
$\mathcal{K}(\beta,x)\equiv\sqrt{\sigma}K(\beta)=\frac{\sin\beta}{\cos^{2}\beta}\left(x+\sqrt{x^{2}-\frac{1}{3\pi}}\right).$ (E.6)
For a physical string to exist, the classical radius $K$ must be real, which requires the discriminant to be non-negative. Thus, the Minkowski version of the Liouville anomaly imposes a rigid chiral existence bound on the dimensionless mass:
$x\geq\frac{1}{\sqrt{3\pi}}.$ (E.7)
Substituting $E$ from Eq. (E.4) back into $\Phi=0$, the mass boundary terms cancel, yielding the parametric angular momentum:
$J(\beta)=\frac{\sigma K^{2}}{4}\left(\beta+\frac{\sin 2\beta}{2}\right)-\frac{1}{6\pi}(\tan\beta-\beta)-\frac{q}{2}.$ (E.8)
To fit the unknown fundamental parameters—the dimensionless constituent quark masses ($x_{u},x_{s},x_{c},x_{b}$) and the universal string tension—to the experimental Particle Data Group (PDG) data, we utilize the action monodromy target $\Phi=0$. Rather than performing an arbitrary least-squares fit on $M^{2}$ or $J$ independently, we mathematically minimize the fundamental physical constraint of the theory itself (see the remark 16.2 for the definition of our effective quark mass).
For any given observed resonance with experimental mass $M$ and spin $J$, we first invert Eq. (E.8) to find the unique theoretical angle $\beta(J)$ such that $J_{\rm param}(\beta)=J$. Substituting this angle back into the master phase equation, the terms corresponding to angular momentum cancel out. Since the theoretical parameters natively satisfy $\Phi=0$, the complicated area, boundary mass, and Liouville anomaly terms collapse, leaving the geometric phase mismatch:
$\Phi(M,J)=2\pi K(\beta)\left(M-E_{\rm param}(\beta)\right).$ (E.9)
This target function has physical and mathematical meaning. Because $K\propto 1/\sqrt{\sigma}$, the target $\Phi$ is naturally dimensionless. At high energies, where $K\propto E/\sigma$, the action monodromy asymptotically scales as $\Phi\propto M\Delta M/\sigma\propto\Delta(M^{2})/(2\sigma)$, intrinsically generating the correct mass-squared weighting expected for Regge trajectories without imposing it ad hoc. Crucially, for the non-rotating ($J=0$) pseudo-Goldstone bosons ($\pi$ and $K$), the string shrinks to a point ($\beta\to 0$ and $K\to 0$), causing $\Phi$ to identically vanish. This naturally decouples the anomalously light chiral masses from the macroscopic string tension.
The global 5D parameter optimization was performed by minimizing the sum of the squared geometric action residuals,
$\chi^{2}(m_{u},m_{s},m_{c},m_{b},\sigma)=\sum_{q_{1},q_{2}}\sum_{J,M}\left(\Phi(M,J,x_{q_{1}q_{2}},\beta(J))\right)^{2},$ (E.10)
across all thirty-six on-shell PDG states spanning the thirteen meson families, where $x_{q_{1}q_{2}}=(x_{q_{1}}+x_{q_{2}})/2$ is the dynamically constructed average boundary mass for the given state.